OpenStudy (anonymous):

use the properties of logarithms to simplify the following f(x) = ln sqrt 10x

OpenStudy (kc_kennylau):

\[\ln a^n=n\ln a\]

OpenStudy (kc_kennylau):

Bear in mind that \(\sqrt m=m^{\frac12}\)

OpenStudy (anonymous):

how do you solve this problem?

OpenStudy (ipwnbunnies):

Use what Kenny just gave you.

OpenStudy (ipwnbunnies):

First step: \[f(x) = \ln(10^{1/2})\]

OpenStudy (kc_kennylau):

Don't forget the x :)

OpenStudy (ipwnbunnies):

T.T \[f(x) = \ln(10^{1/2}x)\]